A 65mm diameter solid shaft is to be welded to a flat plate by a fillet weld
around the circumference of the shaft. Determine the size of weld if the
torque on the shaft is 3KN-m. The allowable shear stress in the weld is 70
N/mm2
1
Expert's answer
2021-09-20T02:48:45-0400


If, h= Size of fillet weld 

t= throat thickness

T= Torque applied 

The maximum shear stress developed in the weld material is given by,

"\tau_s =\frac{2×T}{\pi ×t×d^2}"



Given as 

"\tau_s=66 N/mm^2\\
T=1750 ×1000Nmm\\
d=65mm\\"

Solving for t, we get 

"t=\frac{2T}{\pi×\tau_s×d^2}=\frac{2×1750×1000}{3.14×66×65^2}"


Size of weld(h)="\frac{t}{0.707}= 5.66mm"